# Enhancing Events in Neutrino Telescopes through Deep Learning-Driven Super-Resolution

### v1

@@ II Architecture
Our implementation of super-resolution is organized by a high-level algorithm that compares predicted super-resolved events with their true, unmasked events, as depicted in fig. 2. This is achieved by encoding each event into a 2D image and passing these images through a UNet-type network. The output is also a 2D image representing the super-resolved event. One of the main challenges with this pipeline is adequately encoding and decoding the timing information, which is needed for the prediction of the time series on virtual OMs. This difficulty arises from the geometrically sparse distribution of photon hits for any given neutrino interaction and the highly variable length of the photon hit time series in a given OM. Recently, Ref. [9] demonstrated that an asymmetric Gaussian mixture model can effectively represent the time series of shower-like events in IceCube. Building upon this concept, we train a variational autoencoder (VAE) to encode the time series into a compressed latent space [17]. The UNet then operates fully in the latent space, taking in a point in the latent space for each physical OM and predicting one for each virtual OM. Examples of the VAE representation and super-resolution network predicted time series for two specific virtual OMs are shown in the top plot of fig. 1.
@@ III.1 Virtual Strings
To produce our dataset, we use the open-source neutrino telescope simulation toolkit Prometheus [40]. This allows us to define arbitrary detector geometries. As a benchmark, we use an orthogonal grid of OMs in ice comprising a $15\times 15$ array of strings, each with 61 OMs, for a total of 13,705 OMs. The inter-string spacing is 60 m, and the OMs vertical spacing is 15 m. As input to the network, we mask out every other string, resulting in a $8\times 8$ grid of strings with 120 m spacing. We note that the virtual OM shadow effect, caused by photons being absorbed when hitting virtual OMs in the denser geometry, is negligible.
@@ Appendix A More Details on the UNet
In the particular geometry detailed in the main text, characterized by a dense arrangement of 225 strings with 61 OMs per string, we generate $225\times 61$ images with 68 features at each pixel, as illustrated in fig. 2. The features at each OM contains its 3D coordinates, the number of photon hits, and the 64-parameter latent vector containing the timing information. The inputs to UNet++ are masked, while the unmasked events are retained only for loss computation. As is standard for UNet-type architectures, the outputs are also images of the same spatial size. The output has 65 features at each pixel, corresponding to the number of photon hits and the timing latent vector. MSE loss is used for both the prediction of the number of photon hits and the timing latent vector. The timing latent vector prediction is only trained on OMs that saw light in the event. We use the AdamW optimizer with an initial learning rate of 0.001, with periodic drops when the validation loss plateaus. The network is trained for 30 epochs with a batch size of 128 on an NVIDIA A100 GPU.

### v2

@@ II Architecture
Our implementation of super-resolution is organized by a high-level algorithm that compares predicted super-resolved events with their true, unmasked events, as depicted in fig. 2. This is achieved by encoding each event into a 2D image and passing these images through a UNet-type network. The output is also a 2D image representing the super-resolved event. One of the main challenges with this pipeline is adequately encoding and decoding the timing information, which is needed for the prediction of the time series on virtual OMs. This difficulty arises from the geometrically sparse distribution of photon hits for any given neutrino interaction and the highly variable length of the photon hit time series in a given OM. Recently, Ref. [9] demonstrated that an asymmetric Gaussian mixture model can effectively represent the time series of shower-like events in IceCube. Building upon this concept, we train a variational autoencoder (VAE) to encode the time series into a compressed latent space [17]. The core idea of this process is to take a finely binned distribution of photon hit times (we chose to use 5000 bins, each 1 ns wide) and compress it into a more compact yet effective representation—a 64-parameter latent vector. The UNet then operates fully in the latent space, taking in a point in the latent space for each physical OM and predicting one for each virtual OM. Examples of the VAE representation and super-resolution network predicted time series for two specific virtual OMs are shown in the top plot of fig. 1. While the shapes of most predicted timing PDFs closely match that of the true hits, those farther from the interaction point, with fewer statistics, are harder to reconstruct accurately. Given the limited data on these OMs, it is often sufficient to capture key features like the first hit time and the position of the peak, rather than the exact shape of the distribution.
@@ III.1 Virtual Strings
To produce our dataset, we use the open-source neutrino telescope simulation toolkit Prometheus [40]. This allows us to define arbitrary detector geometries. As a benchmark, we use an orthogonal grid of OMs in ice comprising a $15\times 15$ array of strings, each with 61 OMs, for a total of 13,705 OMs. The inter-string spacing is 60 m, and the OMs vertical spacing is 15 m. As input to the network, we mask out every other string, resulting in a $8\times 8$ grid of strings with 120 m spacing. We note that the virtual OM shadow effect, caused by photons being absorbed when hitting virtual OMs in the denser geometry, is negligible. Thus, in this particular geometry, characterized by a dense arrangement of 225 strings with 61 OMs per string, we generate $225\times 61$ images, as illustrated in fig. 2.
@@ Appendix A More Details on the UNet
The features at each OM contains its 3D coordinates, the number of photon hits, and the 64-parameter latent vector containing the timing information. The inputs to UNet++ are masked, while the unmasked events are retained only for loss computation. As is standard for UNet-type architectures, the outputs are also images of the same spatial size. The output has 65 features at each pixel, corresponding to the number of photon hits and the timing latent vector. MSE loss is used for both the prediction of the number of photon hits and the timing latent vector. The timing latent vector prediction is only trained on OMs that saw light in the event. We use the AdamW optimizer with an initial learning rate of 0.001, with periodic drops when the validation loss plateaus. The network is trained for 30 epochs with a batch size of 128 on an NVIDIA A100 GPU.